OpenStudy (anonymous):

An open box is to be made from a square piece of material, 12 inches on a side, by cutting equal squares from each corner and turning up the sides. Find the dimensions of the box that produces the greatest volume. Answer: 2"by 8" by 8"

OpenStudy (anonymous):

@yun2thejae

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

Eww optimization.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok that was wrong volume is given by \[V(x)=x(12-2x)^2\] that should do it multiply out, take the derivative, set it equal to zero and solve

OpenStudy (anonymous):

@satellite73 volume of a square is \[a^3\] rite?

OpenStudy (anonymous):

so how did u get \[x(12-2x)^3\]

OpenStudy (anonymous):

 ... so we can model the volume as \(V(x)=x(12-2x)^2\). Now we merely need to optimize.\\$$\frac1xV(x)=(12-2x)^2$$If we differentiate, we find:$$-\frac1{x^2}V(x)+\frac1xV'(x)=2(12-2x)(-2)$$If we set \(V'(x)=0\)...$$-\frac1x(12-2x)^2=-4(12-2x)\\-\frac1x(12-2x)=-4\\12-2x=4x\\12=6x\\x=2$$